OpenStudy (anonymous):

2 radical 3 + 6 radical 27 Pleeaasssee heeellppp!!

OpenStudy (anonymous):

2 Radical 3 + 6 Radical 27 = 2 Radical 3 + 18 Radical 3 = 20 Radical 3

OpenStudy (anonymous):

X x X -4X -13=0 HELP PLEASE